OpenStudy (emogirl100):

the volume of the sphere with the radius "r" is v=4/3PIr^3 DIAMETER 16

OpenStudy (emogirl100):

@ShadowLegendX

OpenStudy (wolf1728):

So, you need volume of sphere with diameter of 16?

OpenStudy (emogirl100):

yes

OpenStudy (emogirl100):

but how doi get rid of the diameter

OpenStudy (wolf1728):

radius = diameter divided by 2 radius = 8

OpenStudy (emogirl100):

thx

OpenStudy (emogirl100):

2144.7

OpenStudy (wolf1728):

volume = (4/3) * PI*r^3 volume = (4/3) * PI * 8*8*8 volume = (4/3) * 3.14159 * 512 volume = 2,144.66 Yes that's it !!